OpenStudy (anonymous):

solve the differential equation: dy/dx=(√y)cos^2(√y)

OpenStudy (anonymous):

WHO WILL HELP ME?

OpenStudy (jamesj):

This equation is separable: \[ \int \frac{dy}{\sqrt{y}\ \ \cos^2 \sqrt{y}} = \int dx \] Now substitute \( u = \sqrt{y} \).

OpenStudy (anonymous):

Ummmm but I will face a problem with "cot^2"

OpenStudy (jamesj):

No, with sec^2. And this is the standard derivative of another trig function.

OpenStudy (jamesj):

\[ (d/dx) \tan x = \sec^2 x \]

OpenStudy (anonymous):

Ok.. I will try to solve it.. but just give me "final result"

OpenStudy (anonymous):

My dear!! where are you?

OpenStudy (mr.math):

Are you still facing a problem?

OpenStudy (anonymous):

dy/dx=(√y)cos^2(√y) \[\int\limits_{}^{}dy/\sqrt{y }\cos^2 \sqrt{y}= \int\limits_{}^{}dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\sec^2 \sqrt{y}dy/\sqrt{y}=x+C\]

OpenStudy (anonymous):

\[2\tan \sqrt{y}=x+C\] y=[tan^-1 (x+C)/2]^2

OpenStudy (anonymous):

mark 0 ,,, Thank so much..